# Breather interactions in the integrable discrete Manakov system and trigonometric Yang-Baxter maps

### v1

@@ 1 Introduction
where ${\mathbf{q}}_{n}(t)$ is a 2-component complex vector function of $n\in\mathbb{Z}$, $t\in\mathbb{R}$, and $\sigma=\mp 1$ distinguishes between the focusing/defocusing dispersion regimes. This system, which we will refer to as the integrable discrete Manakov (IDM) system, was introduced in [18, 19, 20] as a vector generalization of the Ablowitz-Ladik model (Eq. (1.1) for a scalar $q_{n}(t)$, see [1, 2]), and it is an integrable, $\mathcal{O}(h^{2})$ spatial discretization of the Manakov system [27]:
@@
where $q(x,t)$ is the 1-soliton solution of the scalar NLS equation, and ${\mathbf{p}}\in\mathbb{C}^{2}$ is a norm-1 complex vector. From a spectral point of view, the physical parameters that characterize the soliton are encoded in: a “discrete eigenvalue” $k=\xi+i\eta\in\mathbb{C}^{+}$, whose real and imaginary parts, respectively, fix the soliton velocity ($v=4\xi$) and amplitude ($A=2\eta>0$); and a “norming constant” ${\mathbf{C}}\in\mathbb{C}^{2}$ associated to the discrete eigenvalue $k$, which determines the center of the soliton, $x_{o}=\ln\sqrt{||{\mathbf{C}}||/2\eta}$, and its “polarization” ${\mathbf{p}}={\mathbf{C}}^{\dagger}/||{\mathbf{C}}||$.
While a 1-soliton solution of the Manakov system is fundamentally governed by the scalar NLS, the vector nature of the solution affects the dynamics when solitons with different polarizations interact. Indeed, as was already established in Manakov’s pioneer paper [27], interacting vector solitons generically change their polarizations upon interaction. Specifically, for a 2-soliton solution with individual solitons traveling at different velocities, in the backward ($t\to-\infty$) and forward ($t\to+\infty$) long-time limits such a solution asymptotically breaks up into individual solitons
@@
[Here and in the following: denotes complex conjugation, and the superscripts T and are used for matrix transpose and conjugate transpose, respectively]. The Manakov formulas (1.4) show that in general when the solitons interact, the intensity distributions of the individual solitons in the two components change, yielding a “polarization shift”, and only when the “initial” polarizations of the solitons ${\mathbf{p}}_{1}^{-},{\mathbf{p}}_{2}^{-}$ are either parallel or orthogonal is the amplitude of each individual component of the solitons conserved. Eqs. (1.4) were obtained in [27] by tracing the asymptotic states of the eigenfunctions through each soliton, multiplying them by the corresponding soliton transmission coefficient, and accounting for bound states of the eigenfunctions via the norming constants. In the multisoliton case, a $J$ soliton collision is equivalent to the composition of $J(J-1)/2$ pairwise interactions taking place in an arbitrary order compatible with the soliton velocities, and the effect of the interactions (i.e., the shifts in the soliton centers and in the polarizations of each individual soliton) is independent of the order in which the pairwise solitons actually interact [5]. In turn, this was shown to be related to the fact that the map $R[k_{1},k_{2}]:({\mathbf{p}}_{1}^{-}\,,\,{\mathbf{p}}_{2}^{-})\mapsto({\mathbf{p}}_{1}^{+}\,,\,{\mathbf{p}}_{2}^{+})$ defined by (1.4) is a (reversible) Yang-Baxter map [5, 14]. Conversely, the matrix refactorization of the transmission coefficients associated to the solitons provides an alternative method to show that the map (1.4) is a Yang-Baxter map [22, 32] [^1] There is a vast literature on Yang-Baxter maps in relation to (fully) discrete integrable systems and consistency around the cube that would be too long to cover here and also not directly relevant to the present work. We refer the interested reader to the book [23]., that is, a solution of the set-theoretical Yang-Baxter equation [17].
@@
The system (1.7) admits a variety of discrete vector soliton solutions, depending on the rank and structure of the norming constant associated with the soliton: fundamental solitons, fundamental breathers (a superposition of two orthogonally polarized fundamental solitons with the same amplitude and velocity, but opposite carrier frequencies), as well as composite breathers (more general superpositions of fundamental solitons). Fundamental solitons are the discrete analog of (1.2), to which they reduce in the continuous limit. On the other hand, fundamental and composite breathers are purely discrete solutions, and do not have a continuous counterpart in the Manakov system. One of the goals of this study is to obtain a complete characterization of soliton-breather and breather-breather interactions for the integrable discrete Manakov system. The formulas for the polarization shifts of discrete fundamental solitons that are the analog of the well-known formulas (1.4) for the interaction of vector solitons in the Manakov system were already obtained in [4, 6, 7]. In this work, we complete the description of the landscape of interactions in the discrete Manakov model by characterizing the interactions between a fundamental soliton and a fundamental breather, between two fundamental breathers, between composite breathers, and between a composite breather and a fundamental soliton or breather. Our study reveals that upon interacting with a fundamental breather, a fundamental soliton becomes a fundamental breather and, conversely, that the interaction of two fundamental breathers generically yields two fundamental breathers with a polarization shifts, but may also result in a fundamental soliton and a fundamental breather. This type of highly non-trivial interaction was discovered for vector solitons of the complex-coupled short pulse equation (ccSPE) in [13], and it is now reported for the first time for a discrete integrable system of NLS-type. Explicit formulas for the coefficients that characterize the fundamental breathers and for their polarization vectors are also obtained. Furthermore, we show that composite breathers interact trivially both with each other and with a fundamental soliton or breather. All interactions are illustrated with plots. The results are then interpreted in terms of a refactorization property for the transmission coefficients associated to each soliton, which produces a novel Yang-Baxter map of trigonometric type.
@@
The structure of the paper is as follows. In Section 2, we give a brief overview of the IST for the IDM system as developed in Ref. [4], and of its 1-soliton and 1-breather solutions. In Section 3, we derive the explicit expressions of the (matrix) transmission coefficients corresponding to a 1-fundamental soliton, a 1-fundamental breather, and a 1-composite breather solution. In Section 4 we perform long-time asymptotics analysis on exact soliton-breather and breather-breather solutions, and we use it in conjunction with the transmission coefficients for a single soliton/breather solution to obtain the maps for the polarization vectors that describe the solitons and breathers in the forward long-time limit $\tau\to+\infty$ in terms of their values as $\tau\to-\infty$. In Section 5 we show, on one hand, how the fundamental breather interaction leads to a Yang-Baxter refactorization property for the transmission coefficients, and, on the other hand, how the map itself can be derived from the refactorization property. Finally, Section 6 is devoted to some concluding remarks, and more technical details are provided in the appendices.
@@ 2 Overview of the IST and soliton/breather solutions
${\mathbf{Q}}_{n}(\tau)=\begin{pmatrix}Q_{n}^{(1)}&Q_{n}^{(2)}\\ (-1)^{n+1}Q_{n}^{(2)\,*}&(-1)^{n}Q_{n}^{(1)\,*}\end{pmatrix},\qquad{\mathbf{R}}_{n}=(-1)^{n}{\mathbf{P}}{\mathbf{Q}}_{n}^{T}{\mathbf{P}},\qquad{\mathbf{P}}=\begin{pmatrix}0&1\\ -1&0\end{pmatrix},$ (2.2)
$z\in\mathbb{C}$ is the spectral parameter, and ${\mathbf{I}}_{2}$ is the $2\times 2$ identity matrix (note that this corresponds to taking $N=M=2$, ${\mathbf{R}}_{n}=-{\mathbf{Q}}_{n}^{\dagger}$, and ${\mathbf{A}}={\mathbf{B}}={\mathbf{I}}_{2}$ in [4]).
@@ 2.1 Direct problem
both approaching the $4\times 4$ identity as $n\to\mp\infty$, respectively. Let $D^{\mp}=\left\{z\in\mathbb{C}:|z|\lessgtr 1\right\}$ denote the interior (--) and the exterior ($+$) of the unit circle $\mathcal{C}=\left\{z\in\mathbb{C}:|z|=1\right\}$. As shown in [4], if the potential ${\mathbf{Q}}_{n}\in\ell^{1}(\mathbb{Z})$ (i.e., with $\sum_{n=-\infty}^{+\infty}\|{\mathbf{Q}}_{n}\|_{a}<\infty$ where $\|{\mathbf{Q}}_{n}\|_{a}$ is any matrix norm of ${\mathbf{Q}}_{n}$), then ${\mathbf{M}}_{n}(z)$, ${\mathbf{N}}_{n}(z)$ defined above are analytic in $D^{+}$ and continuous for $|z|\geq 1$, and $\bar{{\mathbf{M}}}_{n}(z)$, $\bar{{\mathbf{N}}}_{n}(z)$ are analytic for $z\in D^{-}$ and continuous for $|z|\leq 1$. Furthermore, the modified eigenfunctions satisfy the following asymptotics (in their respective regions of analyticity):
@@
where ±$\pm z_{j},\pm\tilde{z}_{j}$ are zeros of $\det{\mathbf{a}}(z)$ (as well as $\det{\mathbf{c}}(z)$) in $D^{+}$, and ±$\pm\bar{z}_{j},\pm\hat{z}_{j}$ are zeros of $\det\bar{{\mathbf{a}}}(z)$ (as well as $\det\bar{{\mathbf{c}}}(z)$) in $D^{-}$. The simplest soliton solutions are obtained assuming that the discrete eigenvalues are simple zeros of $\det{\mathbf{a}}(z)$. But it is possible for a discrete eigenvalue to be a second-order zero of $\det{\mathbf{a}}(z)$ yet still a first-order pole for the meromorphic eigenfunction that appears in the inverse problem. In this respect, both cases should be considered as corresponding to elementary (as opposed to higher order) soliton solutions. Indeed, the following holds. [^2] Analogous results were established in [29] for the matrix NLS equation and in [21] for the ccSPE, and similar arguments can be used here.
@@
Let {±$\left\{\pm z_{j},\pm\bar{z}_{j},\pm\tilde{z}_{j},\pm\hat{z}_{j}\right\}$ be an octet of discrete eigenvalues as in (2.12).
@@
    If ±$\pm z_{j}$ are simple zeros of $\det{\mathbf{a}}(z)$ in $D^{+}$, then $\mathrm{rank}\,{\mathbf{a}}(\pm z_{j})=1$ and ±$\pm z_{j}$ are simple poles for ${\mathbf{M}}_{n}(z){\mathbf{a}}^{-1}(z)$ (and the same holds for the other symmetric eigenvalues in the octet in the respective regions of analyticity).
@@
    If ±$\pm z_{j}$ are double zeros of $\det{\mathbf{a}}(z)$ in $D^{+}$ and ${\mathbf{a}}(\pm z_{j})={\mathbf{0}}_{2}$, then ±$\pm z_{j}$ are still simple poles for ${\mathbf{M}}_{n}(z){\mathbf{a}}^{-1}(z)$ (and the same holds for the other symmetric eigenvalues in the octet in the respective regions of analyticity).
Since in both cases the points ±$\pm z_{j},\pm\tilde{z}_{j}$ (resp., ±$\pm\bar{z}_{j},\pm\hat{z}_{j}$) are simple poles for the function ${\mathbf{M}}{\mathbf{a}}^{-1}$ (resp., $\bar{{\mathbf{M}}}\bar{{\mathbf{a}}}^{-1}$) in $D^{+}$ (resp., $D^{-}$), one can define the corresponding residues as follows:
@@
where ${\mathbf{C}}_{j}$ is the $2\times 2$ norming constant associated to the discrete eigenvalues ±$\pm z_{j}$, and
$\bar{{\mathbf{C}}}_{j}=\bar{z}_{j}^{2}{\mathbf{C}}_{j}^{\dagger},\quad\tilde{{\mathbf{C}}}_{j}=\bar{z}_{j}^{-2}{\mathbf{P}}\bar{{\mathbf{C}}}_{j}^{T}{\mathbf{P}},\quad\hat{{\mathbf{C}}}_{j}=-z_{j}^{-2}{\mathbf{P}}{\mathbf{C}}_{j}^{T}{\mathbf{P}}\,.$ (2.14)
@@ 2.2 Inverse problem
The inverse problem aims at reconstructing the potential ${\mathbf{Q}}_{n}$ in terms of the scattering data (i.e., reflection coefficients, discrete eigenvalues, and associated norming constants). Concretely, one first reconstructs the eigenfunctions from the scattering data, and then the potential is recovered from the asymtptotics of the eigenfunctions in the spectral parameter $z$. The inverse problem for the eigenfunctions is formulated as a Riemann-Hilbert problem (RHP) in the complex variable $z$, for which a suitable normalization condition as $z\to\infty$ must be provided. Since the large-$z$ asymptotic behavior of the modified eigenfunctions depends on the potential ${\mathbf{Q}}_{n}$ (cf. Eqs. (2.7)), it is convenient to pose the problem for the following renormalized matrix functions:
@@ 2.4 Explicit 1-soliton and 1-breather solutions
[…]
@@
$\boldsymbol{\gamma}_{1}^{\perp}={\mathbf{P}}^{T}\boldsymbol{\gamma}_{1}\,,\qquad d_{1}=\log\frac{\|\boldsymbol{\gamma}_{1}\|\|\boldsymbol{\delta}_{1}\|}{\sinh 2a_{1}}\,,$ (2.29b)
and the rest of the parameters are the same as in the fundamental soliton case. The above expression shows that a fundamental breather is a superposition of two orthogonally polarized fundamental solitons, with the same amplitude and velocity, but opposite carrier frequencies. We want to stress that the vector $\boldsymbol{\delta}_{1}$ introduced here is different from the one used in [4] (where $\boldsymbol{\delta}_{1}$ was used to denote the second column of the norming constant), and instead follows the notation introduced in [13] for the ccSPE. Eq. (2.29a) shows that the fundamental breather reduces to a fundamental soliton by setting either $\kappa_{1}=1$ and $\mu_{1}=0$, or $\kappa_{1}=0$ and $\mu_{1}=1$. Of course, one of the two constants $\kappa_{1}$ and $\mu_{1}$ can always be scaled out (as in the solutions presented in [4], where $\mu_{1}=1$ and $\kappa_{1}=\kappa$). However, for the purpose of investigating soliton interactions, it is convenient to keep both constants in. It is worth mentioning that if $\boldsymbol{\gamma}_{1}$ has only one non-zero components, then the opposite carriers are split in different components, and hence the oscillations are suppressed. While with a single breather it is always possible to reduce oneself to such a case by an appropriate axes rotation and exploit the unitary invariance of the system (1.7), this is not possible, in general, with more than one breather.
@@
$\begin{pmatrix}Q_{n}^{(1)}(\tau)\\ Q_{n}^{(2)}(\tau)\end{pmatrix}=\frac{1}{1+g_{n+1}(\tau)}\Big{[}\Theta_{n+1}(\tau)e^{-i(\omega_{1}+ia_{1}v_{1})\tau}\boldsymbol{\gamma}_{1}^{*}+(-1)^{n+1}\Theta_{n+1}^{*}(\tau)e^{i(\omega_{1}-iav_{1})\tau}\boldsymbol{\varepsilon}_{1}^{\perp}\Big{]}\,,$ (2.30a)
@@
$\displaystyle g_{n}(\tau)=\frac{e^{-4a_{1}n+2a_{1}v_{1}\tau}}{\sinh^{2}2a_{1}}\|{\mathbf{C}}_{1}\|^{2}+\Gamma_{n}(\tau)+\Gamma_{n}^{*}(\tau)+s^{4}|\Gamma_{n}(\tau)|^{2}\,,\qquad s=\frac{\cos 2b_{1}}{\sinh 2a_{1}}\,,$ (2.30b)
$\displaystyle\Theta_{n}(\tau)=-2\bar{z}_{1}^{2n}\big{[}1-s^{2}\Gamma_{n}(\tau)\big{]}\,,\;\;\;\Gamma_{n}(\tau)=(-1)^{n}\frac{4}{(\omega_{1}-ia_{1}v_{1})^{2}}z_{1}^{-4n}e^{2i(\omega_{1}-ia_{1}v_{1})\tau}\det{\mathbf{C}}_{1}\,.$ (2.30c)
The above formula for a generic composite breather is novel; only a special case, corresponding to ${\mathbf{C}}_{1}=\left(\boldsymbol{\gamma}_{1}\,,\eta{\mathbf{P}}\,\boldsymbol{\gamma}_{1}^{*}\right)$ with $\eta\in\mathbb{C}$, was given in [6]. We also mention that in [28] soliton solutions for the discrete coupled nonlinear Schrödinger equations (which are essentially the IDM system considered here) are derived using Hirota’s bilinear formalism. However, only fundamental soliton solutions can be obtained from the Pfaffian solutions of the bilinear equations. Some plots of 1-soliton and 1-breather solutions are given in Fig. 1.

### v2

@@ 1 Introduction
where ${\mathbf{q}}_{n}(t)$ is a 2-component complex vector function of $n\in\mathbb{Z}$, $t\in\mathbb{R}$, and $\sigma=\mp 1$ distinguishes between the focusing/defocusing dispersion regimes. This system, which we will refer to as the integrable discrete Manakov (IDM) system, was introduced in [18, 19, 20] as a vector generalization of the Ablowitz-Ladik model (Eq. (1.1) for a scalar $q_{n}(t)$, see [1, 2]), and it is an integrable $\mathcal{O}(h^{2})$ spatial discretization of the Manakov system [27]:
@@
[…]
While a 1-soliton solution of the Manakov system is fundamentally governed by the scalar NLS, the vector nature of the solution affects the dynamics when solitons with different polarizations interact. Indeed, as was already established in Manakov’s pioneering paper [27], interacting vector solitons generically change their polarizations upon interaction. Specifically, for a 2-soliton solution with individual solitons traveling at different velocities, in the backward ($t\to$-$\infty$) and forward ($t\to+\infty$) long-time limits such a solution asymptotically breaks up into individual solitons
@@
[Here and in the following: denotes complex conjugation, and the superscripts T and are used for matrix transpose and conjugate transpose, respectively]. The Manakov formulas (1.4) show that in general when the solitons interact, the intensity distributions of the individual solitons in the two components change, yielding a “polarization shift”, and only when the “initial” polarizations of the solitons ${\mathbf{p}}_{1}^{-},{\mathbf{p}}_{2}^{-}$ are either parallel or orthogonal is the amplitude of each individual component of the solitons conserved. Eqs. (1.4) were obtained in [27] by tracing the asymptotic states of the eigenfunctions through each soliton, multiplying them by the corresponding soliton transmission coefficient, and accounting for bound states of the eigenfunctions via the norming constants. In the multisoliton case, a $J$ soliton collision is equivalent to the composition of J(J1)/2J(J-1)/2 pairwise interactions taking place in an arbitrary order compatible with the soliton velocities, and the effect of the interactions (i.e., the shifts in the soliton centers and in the polarizations of each individual soliton) is independent of the order in which the pairwise solitons actually interact [5]. In turn, this was shown to be related to the fact that the map $R[k_{1},k_{2}]:({\mathbf{p}}_{1}^{-}\,,\,{\mathbf{p}}_{2}^{-})\mapsto({\mathbf{p}}_{1}^{+}\,,\,{\mathbf{p}}_{2}^{+})$ defined by (1.4) is a (reversible) Yang-Baxter map [5, 14]. Conversely, the matrix refactorization of the transmission coefficients associated to the solitons provides an alternative method to show that the map (1.4) is a Yang-Baxter map [22, 32] [^1] There is a vast literature on Yang-Baxter maps in relation to (fully) discrete integrable systems and consistency around the cube that would be too long to cover here and also not directly relevant to the present work. We refer the interested reader to the book [23]., that is, a solution of the set-theoretical Yang-Baxter equation [17].
@@
The system (1.7) admits a variety of discrete vector soliton solutions, depending on the rank and structure of the norming constant associated with the soliton: fundamental solitons, fundamental breathers (a superposition of two orthogonally polarized fundamental solitons with the same amplitude and velocity, but opposite carrier frequencies), as well as composite breathers (more general superpositions of fundamental solitons). Fundamental solitons are the discrete analog of (1.2), to which they reduce in the continuous limit. On the other hand, fundamental and composite breathers are purely discrete solutions, and do not have a continuous counterpart in the Manakov system. One of the goals of this study is to obtain a complete characterization of soliton-breather and breather-breather interactions for the integrable discrete Manakov system. The formulas for the polarization shifts of discrete fundamental solitons that are the analog of the well-known formulas (1.4) for the interaction of vector solitons in the Manakov system were already obtained in [4, 6, 7]. In this work, we complete the description of the landscape of interactions in the discrete Manakov model by characterizing the interactions between a fundamental soliton and a fundamental breather, between two fundamental breathers, between composite breathers, and between a composite breather and a fundamental soliton or breather. Our study reveals that upon interacting with a fundamental breather, a fundamental soliton becomes a fundamental breather and, conversely, that the interaction of two fundamental breathers generically yields two fundamental breathers with polarization shifts, but may also result in a fundamental soliton and a fundamental breather. This type of highly non-trivial interaction was discovered for vector solitons of the complex-coupled short pulse equation (ccSPE) in [13], and it is now reported for the first time for a discrete integrable system of NLS-type. Explicit formulas for the coefficients that characterize the fundamental breathers and for their polarization vectors are also obtained. Furthermore, we show that composite breathers interact trivially both with each other and with a fundamental soliton or breather. All interactions are illustrated with plots. The results are then interpreted in terms of a refactorization property for the transmission coefficients associated to each soliton, which produces a novel Yang-Baxter map of trigonometric type.
@@
The structure of the paper is as follows. In Section 2, we give a brief overview of the IST for the IDM system as developed in Ref. [4], and of its 1-soliton and 1-breather solutions. In Section 3, we derive the explicit expressions of the (matrix) transmission coefficients corresponding to a 1-fundamental soliton, a 1-fundamental breather, and a 1-composite breather solution. In Section 4 we perform long-time asymptotic analysis on exact soliton-breather and breather-breather solutions, and we use it in conjunction with the transmission coefficients for a single soliton/breather solution to obtain the maps for the polarization vectors that describe the solitons and breathers in the forward long-time limit $\tau\to+\infty$ in terms of their values as $\tau\to$-$\infty$. In Section 5 we show, on one hand, how the fundamental breather interaction leads to a Yang-Baxter refactorization property for the transmission coefficients, and, on the other hand, how the map itself can be derived from the refactorization property. Finally, Section 6 is devoted to some concluding remarks, and more technical details are provided in the appendices.
@@ 2 Overview of the IST and soliton/breather solutions
${\mathbf{Q}}_{n}(\tau)=\begin{pmatrix}Q_{n}^{(1)}&Q_{n}^{(2)}\\ (-1)^{n+1}Q_{n}^{(2)\,*}&(-1)^{n}Q_{n}^{(1)\,*}\end{pmatrix},\qquad{\mathbf{R}}_{n}=(-1)^{n+1}\boldsymbol{\sigma}_{2}{\mathbf{Q}}_{n}^{T}\boldsymbol{\sigma}_{2}\,,\qquad\boldsymbol{\sigma}_{2}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$, (2.2)
$z\in\mathbb{C}$ is the spectral parameter, and ${\mathbf{I}}_{2}$ is the $2\times 2$ identity matrix. Note that this corresponds to taking $N=M=2$, ${\mathbf{R}}_{n}=-{\mathbf{Q}}_{n}^{\dagger}$, and ${\mathbf{A}}={\mathbf{B}}={\mathbf{I}}_{2}$ in [4]; and that the matrix $\mathbf{P}$} used throughout [4] is simply $i\boldsymbol{\sigma}_{2}$.
@@ 2.1 Direct problem
both approaching the $4\times 4$ identity as $n\to\mp\infty$, respectively. Let $D^{\mp}=\left\{z\in\mathbb{C}:|z|\lessgtr 1\right\}$ denote the interior (-) and the exterior (++) of the unit circle $\mathcal{C}=\left\{z\in\mathbb{C}:|z|=1\right\}$. As shown in [4], if the potential ${\mathbf{Q}}_{n}\in\ell^{1}(\mathbb{Z})$ (i.e., with $\sum_{n=-\infty}^{+\infty}\|{\mathbf{Q}}_{n}\|_{a}<\infty$ where $\|{\mathbf{Q}}_{n}\|_{a}$ is any matrix norm of ${\mathbf{Q}}_{n}$), then ${\mathbf{M}}_{n}(z)$, ${\mathbf{N}}_{n}(z)$ defined above are analytic in $D^{+}$ and continuous for $|z|\geq 1$, and $\bar{{\mathbf{M}}}_{n}(z)$, $\bar{{\mathbf{N}}}_{n}(z)$ are analytic for $z\in D^{-}$ and continuous for $|z|\leq 1$. Furthermore, the modified eigenfunctions satisfy the following asymptotics (in their respective regions of analyticity):
@@
where $\pm z_{j},\pm\tilde{z}_{j}$ are zeros of $\det{\mathbf{a}}(z)$ (as well as $\det{\mathbf{c}}(z)$) in $D^{+}$, and $\pm\bar{z}_{j},\pm\hat{z}_{j}$ are zeros of $\det\bar{{\mathbf{a}}}(z)$ (as well as $\det\bar{{\mathbf{c}}}(z)$) in $D^{-}$. The simplest soliton solutions are obtained assuming that the discrete eigenvalues are simple zeros of $\det{\mathbf{a}}(z)$. But it is possible for a discrete eigenvalue to be a second-order zero of $\det{\mathbf{a}}(z)$ yet still a first-order pole for the meromorphic eigenfunction that appears in the inverse problem. In this respect, both cases should be considered as corresponding to elementary (as opposed to higher order) soliton solutions. Indeed, the following holds. [^2] Analogous results were established in [29] for the matrix NLS equation and in [21] for the ccSPE, and similar arguments can be used here.
@@
Let $\left\{\pm z_{j},\pm\bar{z}_{j},\pm\tilde{z}_{j},\pm\hat{z}_{j}\right\}$ be an octet of discrete eigenvalues as in (2.12).
@@
    If $\pm z_{j}$ are simple zeros of $\det{\mathbf{a}}(z)$ in $D^{+}$, then $\mathrm{rank}\,{\mathbf{a}}(\pm z_{j})=1$ and $\pm z_{j}$ are simple poles for ${\mathbf{M}}_{n}(z){\mathbf{a}}^{-1}(z)$ (and the same holds for the other symmetric eigenvalues in the octet in the respective regions of analyticity).
@@
    If $\pm z_{j}$ are double zeros of $\det{\mathbf{a}}(z)$ in $D^{+}$ and ${\mathbf{a}}(\pm z_{j})={\mathbf{0}}_{2}$, then $\pm z_{j}$ are still simple poles for ${\mathbf{M}}_{n}(z){\mathbf{a}}^{-1}(z)$ (and the same holds for the other symmetric eigenvalues in the octet in the respective regions of analyticity).
Since in both cases the points $\pm z_{j},\pm\tilde{z}_{j}$ (resp., $\pm\bar{z}_{j},\pm\hat{z}_{j}$) are simple poles for the function ${\mathbf{M}}{\mathbf{a}}^{-1}$ (resp., $\bar{{\mathbf{M}}}\bar{{\mathbf{a}}}^{-1}$) in $D^{+}$ (resp., $D^{-}$), one can define the corresponding residues as follows:
@@
where ${\mathbf{C}}_{j}$ is the $2\times 2$ norming constant associated to the discrete eigenvalues $\pm z_{j}$, and
$\bar{{\mathbf{C}}}_{j}=\bar{z}_{j}^{2}{\mathbf{C}}_{j}^{\dagger},\quad\tilde{{\mathbf{C}}}_{j}=-\bar{z}_{j}^{-2}\boldsymbol{\sigma}_{2}\bar{{\mathbf{C}}}_{j}^{T}\boldsymbol{\sigma}_{2},\quad\hat{{\mathbf{C}}}_{j}=z_{j}^{-2}\boldsymbol{\sigma}_{2}{\mathbf{C}}_{j}^{T}\boldsymbol{\sigma}_{2}\,.$ (2.14)
@@ 2.2 Inverse problem
The inverse problem aims at reconstructing the potential ${\mathbf{Q}}_{n}$ in terms of the scattering data (i.e., reflection coefficients, discrete eigenvalues, and associated norming constants). Concretely, one first reconstructs the eigenfunctions from the scattering data, and then the potential is recovered from the asymptotics of the eigenfunctions in the spectral parameter $z$. The inverse problem for the eigenfunctions is formulated as a Riemann-Hilbert problem (RHP) in the complex variable $z$, for which a suitable normalization condition as $z\to\infty$ must be provided. Since the large-$z$ asymptotic behavior of the modified eigenfunctions depends on the potential ${\mathbf{Q}}_{n}$ (cf. Eqs. (2.7)), it is convenient to pose the problem for the following renormalized matrix functions:
@@ 2.4 Explicit 1-soliton and 1-breather solutions
${\mathbf{C}}_{1}=\left(\boldsymbol{\gamma}_{1}\ {\mathbf{0}}\right),\quad\boldsymbol{\gamma}_{1}\in\mathbb{C}^{2},$ (2.26)
@@
$\boldsymbol{\gamma}_{1}^{\perp}=-i\boldsymbol{\sigma}_{2}\boldsymbol{\gamma}_{1}\,,\qquad d_{1}=\log\frac{\|\boldsymbol{\gamma}_{1}\|\|\boldsymbol{\delta}_{1}\|}{\sinh 2a_{1}}\,$, (2.29b)
and the rest of the parameters are the same as in the fundamental soliton case. The above expression shows that a fundamental breather is a superposition of two orthogonally polarized fundamental solitons, with the same amplitude and velocity, but opposite carrier frequencies. We want to stress that the vector $\boldsymbol{\delta}_{1}$ introduced here is different from the one used in [4] (where $\boldsymbol{\delta}_{1}$ was used to denote the second column of the norming constant), and instead follows the notation introduced in [13] for the ccSPE. Eq. (2.29) shows that the fundamental breather reduces to a fundamental soliton by setting either $\kappa_{1}=1$ and $\mu_{1}=0$, or $\kappa_{1}=0$ and $\mu_{1}=1$. Of course, one of the two constants $\kappa_{1}$ and $\mu_{1}$ can always be scaled out (as in the solutions presented in [4], where $\mu_{1}=1$ and $\kappa_{1}=\kappa$). However, for the purpose of investigating soliton interactions, it is convenient to keep both constants in. It is worth mentioning that if $\boldsymbol{\gamma}_{1}$ has only one non-zero component, then the opposite carriers are split in different components, and hence the oscillations are suppressed. While with a single breather it is always possible to reduce it to such a case by an appropriate axes rotation and exploit the unitary invariance of the system (1.7), this is not possible, in general, with more than one breather.
@@
$\begin{pmatrix}Q_{n}^{(1)}(\tau)\\ Q_{n}^{(2)}(\tau)\end{pmatrix}=\frac{1}{1+g_{n+1}(\tau)}\Big[\Theta_{n+1}(\tau)e^{-2i(\omega_{1}+ia_{1}v_{1})\tau}\boldsymbol{\gamma}_{1}^{*}+(-1)^{n+1}\Theta_{n+1}^{*}(\tau)e^{2i(\omega_{1}-ia_{1}v_{1})\tau}\boldsymbol{\varepsilon}_{1}^{\perp}\Big]\,$, (2.30a)
@@
$\displaystyle g_{n}(\tau)=\frac{e^{-4a_{1}n+{4a_{1}v_{1}\tau}}}{\sinh^{2}2a_{1}}\|{\mathbf{C}}_{1}\|^{2}+\Gamma_{n}(\tau)+\Gamma_{n}^{*}(\tau)+s^{4}|\Gamma_{n}(\tau)|^{2}\,,\qquad s=\frac{\cos 2b_{1}}{\sinh 2a_{1}}\,$, (2.30b)
$\displaystyle\Theta_{n}(\tau)=-2\bar{z}_{1}^{2n}\big[1-s^{2}\Gamma_{n}(\tau)\big]\,,\;\;\;\Gamma_{n}(\tau)=(-1)^{n}\frac{1}{(\omega_{1}-ia_{1}v_{1})^{2}}z_{1}^{-4n}e^{4i(\omega_{1}-ia_{1}v_{1})\tau}\det{\mathbf{C}}_{1}\,.$ (2.30c)
The above formula for a generic composite breather is novel; only a special case, corresponding to ${\mathbf{C}}_{1}=\left(\boldsymbol{\gamma}_{1}\,,i\eta\boldsymbol{\sigma}_{2}\boldsymbol{\gamma}_{1}^{*}\right)$ with $\eta\in\mathbb{C}$, was given in [6]. We also mention that in [28] soliton solutions for the discrete coupled nonlinear Schrödinger equations (which are essentially the IDM system considered here) are derived using Hirota’s bilinear formalism. However, only fundamental soliton solutions can be obtained from the Pfaffian solutions of the bilinear equations that are reported in [28]. Some plots of 1-soliton and 1-breather solutions are given in Fig. 1.
@@
As we will show in Section 4, the key difference between rank-1 and rank-2 solitons is that the former interact nontrivially with each other and with rank-2 solitons. Conversely, when rank-2 solitons interact with each other or with rank-1 solitons, the only effect of the interaction is an overall phase/position shift, while the polarization vectors are left unchanged.